OpenStudy (anonymous):

What is the 35th term of the arithmetic sequence where a1 = 13 and a17 = -83?

OpenStudy (anonymous):

you threre @campbell_st ?

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (campbell_st):

look the formula for a term in an arithmetic series is \[a_{n} = a_{1} + (n -1)\times d\] using the information for the 17th term you have \[-83 = 13 + (17 - 1) \times d\] so solve for d. when you get d, you know the 1st term and n = 35 so substitute the information into the formula to find the 35th term hope it helps

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@ganeshie8 @superhelp101

OpenStudy (campbell_st):

thats the value of d... now go back to the formula with n = 35 and a1 = 13 substitute the value to find the 35th term

OpenStudy (anonymous):

@camerondoherty

OpenStudy (anonymous):

how would you set that up

OpenStudy (anonymous):

@aum

OpenStudy (superhelp101):

well first step is you need to find the common difference..how? we let a17 be our a_n so the formula... An=A1+(n−1)d like i said we let A17 be An that means our n = 17 −51=13+(17−1)d can you try solving d?

OpenStudy (anonymous):

-6

OpenStudy (aum):

\( \large a_{n} = a_{1} + (n -1) * d \\ \large a_n = 13 + (n-1)*(-6) \\ \large a_{35} = 13 + (35-1) * (-6) \)

OpenStudy (campbell_st):

to set it up its direct substitution \[a_{35} = 13 + (35 - 1) \times -6\]

OpenStudy (anonymous):

-191

OpenStudy (aum):

yes.

OpenStudy (anonymous):

What is the 20th term of the arithmetic sequence 3, -3, -9, ... ?

OpenStudy (campbell_st):

ok... so what is the common difference... or what is the pattern saying...?